# Multi-wavelength Study of Dust Emission in the Young Edge-on Protostellar Disk HH 212

### v1

@@ Abstract
Grain growth in disks around young stars plays a crucial role in the formation of planets. Early grain growth has been suggested in the HH 212 protostellar disk by previous polarization observations. To confirm it and to determine the grain size, we analyze high-resolution multi-band observations of the disk obtained with Atacama Large Millimeter/submillimeter Array (ALMA) in Bands 9 (0.4 mm), 7 (0.9 mm), 6 (1.3 mm), 3 (3 mm) as well as with Very Large Array (VLA) in Band Ka (9 mm) and present new VLA data in Bands Q (7 mm), K (1.3 cm), and X (3 cm). We adopt a parameterized flared disk model to fit the continuum maps of the disk in these bands and derive the opacities, albedos, and opacity spectral index $\mathrm{\beta}$ of the dust in the disk, taking into account the dust scattering ignored in the previous work modeling the multi-band data of this source. For the VLA bands, since the continuum emission of the disk is more contaminated by the free-free emission at longer wavelengths, we only include the Band Q data in our modeling. The obtained opacities, albedos, and opacity spectral index $\beta$ (with a value of $\sim$ 1.2) suggest that the upper limit of maximum grain size in the disk be $\sim$ 130 $\mu$m, consistent with that implied in the previous polarization observations in Band 7, supporting the grain growth in this disk.
@@ 1 Introduction
Accurate dust opacities at (sub)millimeter wavelengths are required to estimate the mass of protostellar disks and thus the mass budgets for planet formation. They are also important probes of dust sizes and thus grain growth. From the previous survey at $\mathrm{\lambda}$=1.3mm, the empirical opacity spectral index in T-Tauri disks is estimated to be $\mathrm{\beta=1}$ (Beckwith et al., 1990). A 3mm survey of T-Tauri disks also indicates a similar value of $\mathrm{\beta}$ (Ricci et al., 2010). This infers that the values of $\mathrm{\beta}$ in the Class II T-Tauri disks at millimeter wavelengths are smaller than those in the diffuse interstellar medium (Draine & Lee, 1984; Li & Draine, 2001), suggesting substantial grain growth in the T-Tauri disks for planet formation.
For more embedded Class 0/I stages, (sub)millimeter observations also suggest grain growth for early planet formation. The ALMA/VLA survey of Class 0/I disks in the Perseus and the Orion star-forming regions show a preference for a low value of $\mathrm{\beta\lesssim 1}$ and significant grain growth (Tobin et al., 2020; Tychoniec et al., 2020). Some recent studies adapt certain dust models in radiative transfer modeling of multi-wavelength continuum observations of some Class 0/I disks and infer the maximum grain size of these young sources to be mm/cm (Xu, 2022; Han et al., 2023; Zamponi et al., 2024).
@@
Moreover, recent studies of dust coagulation calculations show that dust grains in the protostellar disk could grow well beyond the fragmentation barrier (in which the growths of dust are suppressed by fragmentation threshold velocity) into the kilogram pebbles at the stage of Class 0/I (Xu & Armitage, 2023), supporting the presence of planet formation in the disks during the Class 0/I phase.
Our target, HH 212, is a well-studied Class 0 protostellar system in Orion ($\sim$400 pc) with a rotating bipolar jet (Lee et al., 2017a). A nearly edge-on disk with an inclination angle of $\sim 87^{\circ}$ relative to the line of sight (Lee et al., 2021b) is seen deeply embedded in an infalling-rotating flattened envelope (Lee et al., 2014, 2017b) that has a centrifugal barrier at $\mathrm{\sim 44}$ au (Lee et al., 2017c). This disk is the first vertically resolved disk at ALMA band 7 showing a cooler dark lane along the major axis (“hamburger-like structure”; Lee et al. 2017b). The HH 212 disk was also observed in the continuum by ALMA at band 9, 6, 3, and VLA band Ka at high angular resolutions. The continuum emission in these bands has been modeled to obtain the dust properties, e.g., opacity spectral index, in the disk (Lin et al., 2021, hereafter Paper 1).
@@
Early grain growth in the HH 212 disk has been suggested in previous polarization observation (Lee et al., 2018). Later work with an opacity spectral index $\mathrm{\beta\sim 1.7}$ (as in the ISM) was able to reproduce the hamburger-like structure of the disk seen at ALMA Band 7, suggesting no grain growth is needed (Galván-Madrid et al., 2018). Paper 1 found that the absorption dust opacity has a spectral index $\mathrm{\beta\sim 1.1}$ similar to the $\mathrm{\beta\sim 1}$ found in T-Tauri disks (Beckwith et al., 1990), but with a lower limit at $\mathrm{\lambda\sim 1.3mm}$ lower than that in T-Tauri disks by a factor of 2, potentially suggesting a somewhat smaller degree of grain growth with different grain properties. Recent well-resolved dust polarization observation at ALMA band 7 revealed the importance of dust self-scattering caused by $\mathrm{\sim 75\mu m}$ grains, suggesting an early grain growth in the HH 212 disk (Lee et al., 2021a).
@@ 2.3 Spectral energy distribution
The $\mathrm{\alpha_{B7-B6}}$ map (Figure 2b) shows a value smaller than 2 within a radius of $\sim$60 au in the midplane, which could be due to radial temperature gradient (Galván-Madrid et al., 2018) or scattering (Zhu et al., 2019; Liu, 2019) in the optically thick disk. From ALMA band 7 and VLA band Q, the value of $\mathrm{\alpha}$ at the center part of the disk tends to gradually increase but is smaller than 2. This suggests that the intensity between ALMA band 7 and VLA band Q is dominated by optically thick thermal emission instead of optically thin emission from large dust grain with small opacity spectral index $\mathrm{\beta}$. From VLA band Q to band X, the low value of $\mathrm{\alpha}<2$ suggests a presence of free-free contamination in the central part of the disk.
To determine the possible contribution from free-free emission, we plot the spectral energy distribution (SED) of the HH 212 disk (Figure 3), in which the total flux of each band is estimated from the region above 3 $\mathrm{\sigma}$. The spectral index $\mathrm{\alpha_{B9-B3}=2.42\pm 0.15}$ derived from the flux densities of ALMA bands shows an optically thick limit in Rayleigh-Jeans tail. For VLA band Q to band K, the $\mathrm{\alpha_{BQ-BK}=2.64\pm 0.43}$ is similar to that from ALMA bands due to a combination of both thermal and free-free emission.
@@ 3.1 Disk model
We assume the pressure scale height $h_{s}$ is proportional to $\frac{c_{s}}{\Omega_{K}}$ within the reference radius, where $c_{s}$ is isothermal sound speed and $\Omega_{K}$ is Keperian frequency. For the outer part of the disk, we assume an exponentially tapered pressure scale height as described in Lee et al. (2017b), which is an empirical necessity and is consistent with the disk formation simulation (Xu & Kunz, 2021).
$h_{s}(R)=\left\{\begin{array}[]{ll}h_{t}\left(\frac{R}{R_{t}}\right)^{1+\frac{1-q}{2}},&R<R_{t},\\ h_{t}\exp\left[\left(-\frac{R-R_{t}}{R_{o}-R_{t}}\right)^{2}\right],&R_{t}<R<R_{o}.\end{array}\right.$ (4)
@@
where $\Sigma_{t}$ is the surface density at the reference radius
@@ 3.2 Dust opacities
We first consider two dust models: DSHARP (Birnstiel et al., 2018) and DIANA (Woitke et al., 2016). DSHARP dust model is composed of water ice (Warren & Brandt, 2008), astronomical silicates (Draine, 2003), refractory organics, and troilite (Henning & Stognienko, 1996) with the adopted volume fractions of 36.4%, 16.7%, 2.6%, and 44.3%, respectively. DIANA dust
Table 3: The summary of different dust opacity assumption.
Dust assumption Parameter Range Best-fit parameters $\tilde{\chi}^{2}$
(1) (2) (3) (4)
DSHARP dust model 20$\mathrm{<Rt(au)<}$40 $\mathrm{R_{t}=40^{+2}_{-0}}$ $\mathrm{\tilde{\chi}_{B9}^{2}=9.284}$
(Birnstiel et al. 2018) 60$\mathrm{<Tt(K)<}$120 $\mathrm{T_{t}=102\pm 6}$ $\mathrm{\tilde{\chi}_{B7}^{2}=6.064}$
$\mathrm{10^{1}<a_{max}(\mu m)<10^{3}}$ $\mathrm{a_{max}=130\pm 20}$ $\mathrm{\tilde{\chi}_{B6}^{2}=4.117}$
1.0$\mathrm{<Q_{t}<}$2.5 $\mathrm{Q_{t}=1.0^{+0.1}_{-0.0}}$ $\mathrm{\tilde{\chi}_{B3}^{2}=4.140}$
$\mathrm{\tilde{\chi}_{BQ}^{2}=5.859}$
DIANA dust model 20$\mathrm{<Rt(au)<}$40 $\mathrm{R_{t}=36\pm 2}$ $\mathrm{\tilde{\chi}_{B9}^{2}=9.266}$
(Woitke et al. 2016) 60$\mathrm{<Tt(K)<}$120 $\mathrm{T_{t}=60^{+6}_{-0}}$ $\mathrm{\tilde{\chi}_{B7}^{2}=6.069}$
$\mathrm{10^{1}<a_{max}(\mu m)<10^{3}}$ $\mathrm{a_{max}=50\pm 20}$ $\mathrm{\tilde{\chi}_{B6}^{2}=4.129}$
1.0$\mathrm{<Q_{t}<}$2.5 $\mathrm{Q_{t}=2.0\pm 0.1}$ $\mathrm{\tilde{\chi}_{B3}^{2}=4.138}$
$\mathrm{\tilde{\chi}_{BQ}^{2}=5.859}$
Parameterized dust 20$\mathrm{<Rt(au)<}$40 $\mathrm{R_{t}=30\pm 2}$ $\mathrm{\kappa_{B9,abs}=1.76\pm 0.13}$ $\mathrm{\omega_{B9,eff}=0.50\pm 0.05}$ $\mathrm{\tilde{\chi}_{B9}^{2}=}$9.265
opacity (PDO) model 60$\mathrm{<Tt(K)<}$120 $\mathrm{T_{t}=72\pm 6}$ $\mathrm{\kappa_{B7,abs}=1.61\pm 0.15}$ $\mathrm{\omega_{B7,eff}=0.30\pm 0.05}$ $\mathrm{\tilde{\chi}_{B7}^{2}=}$6.061
($\mathrm{Q_{t}}$=1) $\mathrm{10^{-1.5}<\kappa_{\nu}(cm^{2}/g)<10^{0.5}}$ $\mathrm{\kappa_{B6,abs}=1.33\pm 0.19}$ $\mathrm{\omega_{B6,eff}=0.10\pm 0.05}$ $\mathrm{\tilde{\chi}_{B6}^{2}=}$4.112
$\mathrm{0.0<\omega_{\nu,eff}<1.0}$ $\mathrm{\kappa_{B3,abs}=0.47\pm 0.06}$ $\mathrm{\omega_{B3,eff}=0.10\pm 0.05}$ $\mathrm{\tilde{\chi}_{B3}^{2}=}$4.133
$\mathrm{\kappa_{BQ,abs}=0.14\pm 0.01}$ $\mathrm{\omega_{BQ,eff}=0.00^{+0.05}_{-0.00}}$ $\mathrm{\tilde{\chi}_{BQ}^{2}=}$5.858
model consists of pyroxene with 70% Magnesium and carbon. Assuming spherical compact dust grains with MRN grain size distribution (Mathis et al., 1977, n(a)$\mathrm{\propto a^{-3.5}}$) and a maximum grain size $\mathrm{a_{max}}$, we use OPTOOL (Dominik et al., 2021) to generate the corresponding dust absorption and scattering opacities for radiative transfer (Figure 4). The grain size at different heights of the disk may be different. In this study, we assume a constant $\mathrm{a_{max}}$ throughout the disk for simplicity.
@@
where $\mathrm{\kappa_{\nu,abs}}$, $\mathrm{\kappa_{\nu,sca}}$ and $\mathrm{g_{\nu}}$ are absorption opacity, scattering opacity and anisotropic scattering factor (D’Alessio et al., 2001). Due to different compositions, the absorption opacities of the DIANA dust model are about one order larger than those of the DSHARP dust model at (sub)millimeter wavelengths. We set $\mathrm{R_{t}}$, $\mathrm{T_{t}}$, $\mathrm{a_{max}}$, and $\mathrm{Q_{t}}$ as free parameters for these two dust models, which correspond to the geometry, temperature profile, opacity, and density profile of the disk respectively. Because of the uncertain dust properties in Class 0/I disks, we also introduce a Parameterized dust opacity (PDO) model that is not constrained by a specific dust composition. In this case, we assume $\mathrm{Q_{t}=1}$ (similar as Paper 1) and set $\mathrm{R_{t}}$, $\mathrm{T_{t}}$, $\mathrm{\kappa_{\nu,abs}}$, and $\mathrm{\omega_{\nu,eff}}$ as free parameters in each band.
@@ 4.1 Model fitting results
We obtain the best-fit parameters for each dust opacity assumption with the smallest reduced chi-square $\mathrm{\tilde{\chi}^{2}}$ of ALMA and VLA bands. Because of the free-free contamination in centimeter wavelengths (Section 2.3), we only use the ALMA bands 9, 7, 6, 3, and VLA band Q observations to find the best-fit parameters.
Table 3 shows the best-fit parameters and reduced chi-square $\mathrm{\tilde{\chi}^{2}}$ of each dust model. The reduced chi-square $\mathrm{\tilde{\chi}^{2}}$ of ALMA band 9 in each dust model is larger than other bands because of higher contamination from the envelope. The reduced chi-square $\mathrm{\tilde{\chi}^{2}}$ of the PDO model shows a better fitting result than the other two models. This is because the PDO model directly sets opacity in each wavelength as a free parameter, while other dust models’ opacities are constrained by maximum grain size. Therefore, the PDO model is more flexible and can fit the observations better than the DSHARP and DIANA dust models.

### v2

# Early Grain Growth in the Young Protostellar Disk HH 212 Supported by Dust Self-Scattering Modeling

@@ Abstract
Grain growth in disks around young stars plays a crucial role in the formation of planets. Early grain growth has been suggested in the HH 212 protostellar disk by previous polarization observations. To confirm it and to determine the grain size, we analyze high-resolution multi-band observations of the disk obtained with Atacama Large Millimeter/submillimeter Array (ALMA) in Bands 9 (0.4 mm), 7 (0.9 mm), 6 (1.3 mm), 3 (3 mm) as well as with Very Large Array (VLA) in Band Ka (9 mm) and present new VLA data in Bands Q (7 mm), K (1.3 cm), and X (3 cm). We adopt a parameterized flared disk model to fit the continuum maps of the disk in these bands and derive the opacities, albedos, and opacity spectral index $\mathrm{\beta}$ of the dust in the disk, taking into account the dust scattering ignored in the previous work modeling the multi-band data of this source. For the VLA bands, we only include the Band Q data in our modeling to avoid free-free emission contamination. The obtained opacities, albedos, and opacity spectral index $\beta$ (with a value of $\sim$ 1.2) suggest that the upper limit of maximum grain size in the disk be $\sim$ 130 $\mu$m, consistent with that implied in the previous polarization observations in Band 7, supporting the grain growth in this disk. The values of the absorption opacities further highlight the need for a new dust composition model for Class 0/I disks.
@@ 1 Introduction
Accurate dust opacities at (sub)millimeter wavelengths are required to estimate the mass in solids of protostellar disks and thus the mass budgets for planet formation. They are also important probes of dust sizes and thus grain growth. From the previous survey at $\mathrm{\lambda}$=1.3mm, the empirical opacity spectral index in T-Tauri disks is estimated to be $\mathrm{\beta=1}$ (Beckwith et al., 1990). A 3mm survey of T-Tauri disks also indicates a similar value of $\mathrm{\beta}$ (Ricci et al., 2010). This infers that the values of $\mathrm{\beta}$ in the Class II T-Tauri disks at millimeter wavelengths are smaller than those in the diffuse interstellar medium (Draine & Lee, 1984; Li & Draine, 2001), suggesting substantial grain growth in the T-Tauri disks for planet formation.
For more embedded Class 0/I stages, (sub)millimeter observations also suggest grain growth for early planet formation. The ALMA/VLA survey of Class 0/I disks in the Perseus and the Orion star-forming regions show a preference for a low value of $\mathrm{\beta\lesssim 1}$ and significant grain growth (Tobin et al., 2020; Tychoniec et al., 2020). Some recent studies adopt certain dust models in radiative transfer modeling of multi-wavelength continuum observations of some Class 0/I disks and infer the maximum grain size of these young sources to be mm/cm (Xu, 2022; Han et al., 2023; Zamponi et al., 2024).
@@
Moreover, recent studies of dust coagulation calculations show that dust grains in the protostellar disk could grow well beyond the fragmentation barrier (in which the growths of dust are suppressed by fragmentation threshold velocity) into the cm-sized pebbles at the stage of Class 0/I (Xu & Armitage, 2023), supporting the presence of planet formation in the disks during the Class 0/I phase.
Our target, HH 212, is a well-studied Class 0 protostellar system in Orion ($\sim$400 pc) with a rotating bipolar jet (Lee et al., 2017a). A nearly edge-on disk with an inclination angle of $\sim 87^{\circ}$ relative to the line of sight (Lee et al., 2021b) is seen deeply embedded in an infalling-rotating flattened envelope (Lee et al., 2014, 2017b) that has a centrifugal barrier at $\mathrm{\sim 44}$ au (Lee et al., 2017c). This disk is the first vertically resolved disk at ALMA band 7 showing a cooler dark lane along the major axis (“hamburger-like structure”; Lee et al. 2017b). The HH 212 disk was also observed in the continuum by ALMA at band 9, 6, 3, and VLA band Ka at high angular resolutions. Although face-on disks are generally easier to constrain the disk model, specifically for the density and temperature distribution in the radial direction, as compared to edge-on disks. The multi-band observations can also constrain the radial distribution of density and temperature in the edge-on disk HH 212, as different wavelengths probe different radii. The continuum emission in ALMA band 7, 6 and 3 has been modeled to obtain the dust properties, e.g., opacity spectral index, in the disk (Lin et al., 2021, hereafter Paper 1).
@@
Early grain growth in the HH 212 disk has been suggested in previous polarization observation (Lee et al., 2018). Later work with an opacity spectral index $\mathrm{\beta\sim 1.7}$ (as in the ISM) was able to reproduce the hamburger-like structure of the disk seen at ALMA Band 7, suggesting no grain growth is needed (Galván-Madrid et al., 2018). Paper 1 found that the opacity spectral index $\mathrm{\beta\sim 1.1}$ is similar to the $\mathrm{\beta\sim 1}$ found in T-Tauri disks (Beckwith et al., 1990). However, the resulting absorption opacity at $\mathrm{\lambda\sim}$ 1.3mm in Paper 1 is lower than that in T-Tauri disks by a factor of 2, potentially suggesting a somewhat smaller degree of grain growth or different grain properties or both. Recent well-resolved dust polarization observation at ALMA band 7 revealed the importance of dust self-scattering caused by $\mathrm{\sim 75\mu m}$ grains, suggesting an early grain growth in the HH 212 disk (Lee et al., 2021a).
@@ 2.3 Spectral energy distribution
The $\mathrm{\alpha_{B7-B6}}$ map (Figure 2b) shows a value smaller than 2 within a radius of $\sim$60 au in the midplane, which could be due to radial temperature gradient (Galván-Madrid et al., 2018) or scattering (Zhu et al., 2019; Liu, 2019) in the optically thick disk. From ALMA band 7 to VLA band Q, the value of $\mathrm{\alpha}$ at the center part of the disk tends to gradually increase but is smaller than 2. In addition, previous works already showed that the center of this disk is optically thick at ALMA bands 7, 6, and 3 (Lee et al. 2017b; Paper 1). At longer wavelengths, the thermal emission becomes optically thick in the inner region, where the temperature is higher. This will result in a spectral index smaller than 2. This suggests that the intensity between ALMA band 7 and VLA band Q is dominated by optically thick thermal emission instead of optically thin emission from large dust grain with small opacity spectral index $\mathrm{\beta}$. From VLA band Q to band X, the low value of $\mathrm{\alpha}<2$ suggests a presence of free-free contamination in the central part of the disk.
To determine the possible contribution from free-free emission, we plot the spectral energy distribution (SED) of the HH 212 disk (Figure 3), in which the total flux of each band is estimated from the region above 3 $\mathrm{\sigma}$. The spectral index $\mathrm{\alpha_{B9-B3}=2.42\pm 0.15}$ derived from the flux densities of ALMA bands shows that this disk is partially optically thick, with optically thick and thin thermal emission from the inner and outer region of the disk, respectively. For VLA band Q to band K, the $\mathrm{\alpha_{BQ-BK}=2.64\pm 0.43}$ is similar to that from ALMA bands due to a combination of both thermal and free-free emission.
@@ 3.1 Disk model
We assume the pressure scale height $h_{s}$ is proportional to $\frac{c_{s}}{\Omega_{K}}$ within the reference radius, where $c_{s}$ is isothermal sound speed and $\Omega_{K}$ is Keperian frequency. For the outer part of the disk, to reproduce the thinner disk height in the observations, we assume an exponentially tapered pressure scale height as described in Lee et al. (2017b), which is an empirical necessity and is consistent with the disk formation simulation (Xu & Kunz, 2021). This tapered pressure height is likely due to optical depth effects or self-shielding in the disk (Dullemond & Dominik, 2004).
$h_{s}(R)=\left\{\begin{array}[]{ll}h_{t}\left(\frac{R}{R_{t}}\right)^{1+\frac{1-q}{2}},&R<R_{t}\\ h_{t}\exp\left[\left(-\frac{R-R_{t}}{R_{o}-R_{t}}\right)^{2}\right],&R_{t}<R<R_{o}\end{array}\right.$ (4)
@@
where $\Sigma_{t}$ is the surface density at the reference radius. We calculate $\Sigma_{t}$ of the disk by integrating up to $\mathrm{h_{o}}$ because the regions above $\mathrm{h_{o}}$ are likely excavated by the outflow (Lee et al., 2017c) and the disk wind (Lee et al., 2021b).
@@ 3.2 Dust opacities
We first consider two dust models: DSHARP (Birnstiel et al., 2018) and DIANA (Woitke et al., 2016). DSHARP dust model is composed of water ice (Warren & Brandt, 2008), astronomical silicates (Draine, 2003), refractory organics, and troilite (Henning & Stognienko, 1996) with the adopted volume fractions of 36.4%, 16.7%, 2.6%, and 44.3%, respectively. DIANA dust model consists of pyroxene with 70% Magnesium and carbon. Assuming spherical compact dust grains with MRN grain size distribution (Mathis et al., 1977, n(a)$\mathrm{\propto a^{-3.5}}$) and a maximum grain size $\mathrm{a_{max}}$, we use OPTOOL (Dominik et al., 2021) to generate the corresponding dust absorption and scattering opacities for radiative transfer (Figure 4). The grain size at different heights of the disk may be different. In this study, we assume a constant $\mathrm{a_{max}}$ throughout the disk for simplicity.
@@
where $\mathrm{\kappa_{\nu,abs}}$, $\mathrm{\kappa_{\nu,sca}}$ and $\mathrm{g_{\nu}}$ are absorption opacity, scattering opacity and anisotropic scattering factor (D’Alessio et al., 2001). Due to different compositions, the absorption opacities of the DIANA dust model are about one order of magnitude larger than those of the DSHARP dust model at (sub)millimeter wavelengths. We set $\mathrm{R_{t}}$, $\mathrm{T_{t}}$, $\mathrm{a_{max}}$, and $\mathrm{Q_{t}}$ as free parameters for these two dust models, which correspond to the geometry, temperature profile, opacity, and density profile of the disk respectively. Because of the uncertain dust properties in Class 0/I disks, we also introduce a Parameterized dust opacity (PDO) model that is not constrained by a specific dust composition. In this case, we assume $\mathrm{Q_{t}=1}$ (similar as Paper 1) and set $\mathrm{R_{t}}$, $\mathrm{T_{t}}$, $\mathrm{\kappa_{\nu,abs}}$, and $\mathrm{\omega_{\nu,eff}}$ as free parameters in each band.
@@ 4.1 Model fitting results
We obtain the best-fit parameters for each dust opacity assumption with the smallest reduced chi-square $\mathrm{\tilde{\chi}^{2}}$
Table 3: The summary of different dust opacity assumption.
Dust assumption Parameter Range Best-fit parameters $\tilde{\chi}^{2}$
(1) (2) (3) (4)
DSHARP dust model 20$\mathrm{<Rt(au)<}$40 $\mathrm{R_{t}=40^{+2}_{-0}}$ $\mathrm{\tilde{\chi}_{B9}^{2}=19.21}$
(Birnstiel et al. 2018) 60$\mathrm{<Tt(K)<}$120 $\mathrm{T_{t}=102\pm 6}$ $\mathrm{\tilde{\chi}_{B7}^{2}=7.95}$
$\mathrm{10^{1}<a_{max}(\mu m)<10^{3}}$ $\mathrm{a_{max}=130\pm 20}$ $\mathrm{\tilde{\chi}_{B6}^{2}=6.01}$
1.0$\mathrm{<Q_{t}<}$2.5 $\mathrm{Q_{t}=1.0^{+0.1}_{-0.0}}$ $\mathrm{\tilde{\chi}_{B3}^{2}=7.39}$
$\mathrm{\tilde{\chi}_{BQ}^{2}=2.31}$
DIANA dust model 20$\mathrm{<Rt(au)<}$40 $\mathrm{R_{t}=36\pm 2}$ $\mathrm{\tilde{\chi}_{B9}^{2}=16.46}$
(Woitke et al. 2016) 60$\mathrm{<Tt(K)<}$120 $\mathrm{T_{t}=60^{+6}_{-0}}$ $\mathrm{\tilde{\chi}_{B7}^{2}=8.44}$
$\mathrm{10^{1}<a_{max}(\mu m)<10^{3}}$ $\mathrm{a_{max}=50\pm 20}$ $\mathrm{\tilde{\chi}_{B6}^{2}=11.88}$
1.0$\mathrm{<Q_{t}<}$2.5 $\mathrm{Q_{t}=2.0\pm 0.1}$ $\mathrm{\tilde{\chi}_{B3}^{2}=1.84}$
$\mathrm{\tilde{\chi}_{BQ}^{2}=2.17}$
Parameterized dust 20$\mathrm{<Rt(au)<}$40 $\mathrm{R_{t}=30\pm 2}$ $\mathrm{\kappa_{B9,abs}=1.76\pm 0.13}$ $\mathrm{\omega_{B9,eff}=0.50\pm 0.05}$ $\mathrm{\tilde{\chi}_{B9}^{2}=9.10}$
opacity (PDO) model 60$\mathrm{<Tt(K)<}$120 $\mathrm{T_{t}=72\pm 6}$ $\mathrm{\kappa_{B7,abs}=1.61\pm 0.15}$ $\mathrm{\omega_{B7,eff}=0.30\pm 0.05}$ $\mathrm{\tilde{\chi}_{B7}^{2}=2.03}$
($\mathrm{Q_{t}}$=1) $\mathrm{10^{-1.5}<\kappa_{\nu}(cm^{2}/g)<10^{0.5}}$ $\mathrm{\kappa_{B6,abs}=1.33\pm 0.19}$ $\mathrm{\omega_{B6,eff}=0.10\pm 0.05}$ $\mathrm{\tilde{\chi}_{B6}^{2}=6.00}$
$\mathrm{0.0<\omega_{\nu,eff}<1.0}$ $\mathrm{\kappa_{B3,abs}=0.47\pm 0.06}$ $\mathrm{\omega_{B3,eff}=0.10\pm 0.05}$ $\mathrm{\tilde{\chi}_{B3}^{2}=1.43}$
$\mathrm{\kappa_{BQ,abs}=0.14\pm 0.01}$ $\mathrm{\omega_{BQ,eff}=0.00^{+0.05}_{-0.00}}$ $\mathrm{\tilde{\chi}_{BQ}^{2}=1.27}$
of ALMA and VLA bands. Because of the free-free contamination in centimeter wavelengths (Section 2.3), we only use the ALMA bands 9, 7, 6, 3, and VLA band Q observations to find the best-fit parameters.
Table 3 shows the best-fit parameters and reduced chi-square $\mathrm{\tilde{\chi}^{2}}$ of each dust model. The reduced chi-square $\mathrm{\tilde{\chi}^{2}}$ of ALMA band 9 in each dust model is larger than other bands because of higher contamination from the envelope. The reduced chi-square $\mathrm{\tilde{\chi}^{2}}$ of the PDO model shows a better fitting result than the other two models. This is because the PDO model directly sets opacity in each wavelength as a free parameter, while other dust models’ opacities are constrained by maximum grain size. Therefore, the PDO model is more flexible and can fit the observations better than the DSHARP and DIANA dust models. The absorption opacities and effective albedos inferred by the PDO model from ALMA Band 7 to VLA Band Q fall between those predicted by the DSHARP and DIANA models (Figures 8 and 9). A detailed discussion of these results will follow in Section 5.